OpenStudy (anonymous):

If tanA=2/3 and sinB =5/√41 and angles A and B are in Quadrant I, find the value of tan (A+B) showing all steps

OpenStudy (dumbcow):

use angle sum formula for tan \[\tan(A+B) =\frac{tanA + tanB}{1-tanAtanB}\] from sinB = 5/sqrt41 use trig identity sin^2+cos^2 =1 cosB = sqrt(1 - sinB^2) cosB = 4/sqrt41 tanB = sinB/cosB = 5/4 tanA = 2/3 substitute these values into above formula tan(A+B) = 23/2